OpenStudy (anonymous):

PLEASE HELP? What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

consider the eq 2x+4= 8 so solving it we find 2x=8-4 i.e. 2x=4 i.e. x=4/2 x=2 now if we sunstitute x=2 in the eq we must find LHS= RHS Let us take LHS=2x+4 =2*2+4=4+4=8 =RHS thus x=2 is the solution of the given eq.